OpenStudy (anonymous):

how do i solve for x in the problem 0.13x+0.03(60,000-x)=5,000

OpenStudy (anonymous):

first multiply out the 0.03

OpenStudy (anonymous):

then keep anything with an x on it on one side, and all the other ones on another side

OpenStudy (anonymous):

then combine the x's so that you get one coefficient and one x, then divide the otherside by your coefficient and you have your x value

OpenStudy (anonymous):

make sure you get 32000 when you are done

OpenStudy (anonymous):

.13x+0.03(60000-x)=5000 .13x+1800-.03x=5000 .13x-.03x=5000-1800 .13x-.03x=3200 .10x=3200 x=3200/.1 x=32000